find the average rate of change for the following function.

f(x)=2x^3-5x^2+7 between x=-2 and x=1.

The average rate of change for f(x) over the interval -2 to 1 is ___. (Type an integer or a simplified fraction.)
Answer #1

Sol:

average rate of change for the following function.is

=f(1)-f(-2)/(1-(-2)

f(1)=2*(1)^3-5*1^2+7=2-5+7=4

f(-2)=2*(-2)^3-5*(-2)^2+7=-29

=4-(-29)/1-([-2)

=4+29/1+2

=33/3

=11

The average rate of change for f(x) over the interval -2 to 1 is 11